y-2=1/3(x+1) what is the y-intercept form and what does is look like on a graph?
Answer by mananth(16946)

y-2=1/3(x+1)
you have to rewrite the equation in the form y = mx+b
where (x,y) are the co-ordinates and m the slope, b the y intercept
multiply by 3
3(y-2) x+1
3y-6 = x+1
3y= x+7
/3
y = (1/3)x+7
This is the equation in slope intercept form

m.ananth@hotmail.ca